What is another point on the line (-5,86) and (95,16)?
Answer by amarjeeth123(569)

Let the two points be A(-5,86) and B(95,16).
Imagine a straight line between A and B.
The midpoint of these two points will also lie on the same straight line.
Midpoint of AB=C
X-coordinate of C=(-5+95)/2=90/2=45
Y-coordinate of C=(86+16)/2=102/2=51
C(45,51) also lies on the same straight line.